Margin of Safety

Chillmax Company plans to sell 3,500 pairs of shoes at $60 each in the coming year. Unit variable cost is $21 (includes direct materials, direct labor, variable factory overhead, and variable selling expense). Total fixed cost equals $78,000 (includes fixed factory overhead and fixed selling and administrative expense). Break-even units equal 2,000.

Required:

1. Calculate the margin of safety in terms of the number of units.

units

2. Calculate the margin of safety in terms of sales revenue.
Answer #1
1) margin of safety in terms of no of unit
Margin of safety = Total sale unit-breakeven sale unit
Margin of safety = 3500-2000 = 1500 Units
2) margin of safety in term of sale revenue
margin of safety = 1500*60 = 90000

Please give a thumbs up if it is helpful & let me know if any doubt